# MatDecompSDF: High-Fidelity 3D Shape and PBR Material Decomposition from Multi-View Images

### v1

@@ 1. Introduction
The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry (Guo et al., 2023). Traditionally, creating such photorealistic digital assets is a labor-intensive process, often requiring expert artists to manually model geometry and define complex material properties. The advent of neural scene representations has marked a paradigm shift, offering a new avenue for synthesizing and reconstructing the 3D world directly from images. Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (Mildenhall et al., 2020). This seminal work spurred a wave of research exploring dynamic scenes (Pumarola et al., 2021; Li et al., 2021b) and even articulated subjects like humans (Liu et al., 2021; Li et al., 2023; Wu et al., 2023). However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
@@
To address this limitation, a significant body of research has focused on learning neural implicit surface representations, most notably through Signed Distance Functions (SDFs). Frameworks like NeuS (Wang et al., 2021) and VolSDF (Yariv et al., 2021) ingeniously bridge the gap between volume rendering and surface rendering, enabling the extraction of detailed and coherent meshes from multi-view images. The pursuit of efficiency and detail in these reconstructions continues, with methods exploring explicit representations like tensor decompositions (Jin et al., 2022; Fridovich-Keil et al., 2023) and optimization ”baking” strategies to accelerate rendering (Chen et al., 2023a). This hierarchical refinement of detail, from coarse volumetric fields to precise surfaces, mirrors the effectiveness of hierarchical models in other domains, such as understanding the nested structure of language in sign language translation (Guo et al., 2018, 2019b) or parsing complex scenes in visual dialog (Guo et al., 2020, 2021).
@@
Nevertheless, geometry alone is insufficient for true photorealism. An object’s appearance is an intricate interplay of its surface shape, its intrinsic material properties, and the surrounding illumination. The ultimate goal is thus inverse rendering: the challenging, ill-posed problem of disentangling these coupled factors from 2D observations. Early explorations in the neural domain sought to factorize NeRF into reflectance and illumination components (Zhou et al., 2021; Boss et al., 2021), often under simplified assumptions. Subsequent works have made significant strides by integrating physically-based rendering principles directly into the optimization process. These methods aim to decompose appearance into standardized Physically-Based Rendering (PBR) parameters—such as albedo, roughness, and metallic—by leveraging differentiable renderers (Zhang et al., 2021b, 2022b). This disentanglement is critical, as failing to separate these components can lead to ambiguity, a challenge analogous to disambiguating events from correlated signals in audio-visual analysis (Zhou et al., 2022) or grounding language in proposal-free video understanding (Li et al., 2021a).
@@
State-of-the-art inverse rendering frameworks have achieved remarkable results, capable of extracting high-quality triangular meshes along with their PBR material maps and complex lighting environments (Munkberg et al., 2022; Zhang et al., 2023). These methods have been extended to handle complex indoor scenes (Yen-Chen et al., 2021; Chang et al., 2023), unstructured ”in-the-wild” online image collections (Liu et al., 2022), and even animatable human avatars with relightable properties (Zhao et al., 2022). The development of comprehensive frameworks has further unified these disparate efforts (Huang et al., 2023b). Despite this progress, significant challenges remain. First, many high-fidelity methods are computationally expensive, hindering their practical application; recent advances in real-time relighting with 3D Gaussians (Huang et al., 2024c) and efficient representations (Chen et al., 2022) are beginning to address this. The need for efficient models is a universal concern, echoing the importance of model compression techniques like multi-objective convex quantization in deploying large networks (Fan et al., 2024). Second, most models are limited to simple, opaque materials, struggling with optically complex phenomena like translucency (Yu et al., 2023), anisotropic reflectance (Shen et al., 2023), or the intricate structure of hair (Chen et al., 2023b). Third, robustly handling diverse real-world data requires sophisticated techniques, from modeling visibility and occlusion (Zhang et al., 2021a; Srinivasan et al., 2023) to leveraging specialized information like polarization (Mochida et al., 2022).
@@
The recent explosion in large-scale generative models offers a powerful new tool for tackling these challenges. Text-to-2D diffusion models have been ingeniously repurposed as powerful 3D priors through score distillation sampling (Poole et al., 2023), enabling text-to-3D content creation (Lin et al., 2023). These generative techniques are evolving to produce higher-fidelity and more diverse results (Wang et al., 2023) and are increasingly capable of disentangling geometry and appearance in a text-guided manner (Chen et al., 2023c). This allows for semantic texturing (Richardson et al., 2023) and material assignment (Seo et al., 2023), where textual concepts like ”metallic” or ”wooden” are translated into PBR values, a process informed by material-aware image generation (Xu et al., 2023). Furthermore, multi-modal models like CLIP facilitate direct, text-driven editing of neural scenes (Wang et al., 2022), enabling intuitive control over object materials (Dadashnialehi et al., 2023) and textures (Baatz et al., 2022). This convergence of vision, language, and graphics points towards a future of highly controllable and semantic 3D content creation, akin to how cross-modal understanding enhances tasks like sign language production (Tang et al., 2022) and crowd counting (Guo et al., 2019a).
@@
Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals. The ability to recognize human activities (Huang et al., 2020, 2023a), gestures (Gu et al., 2023a), emotions (Gu et al., 2023b), and even vital pulmonary functions (Zhao et al., 2025; Wang et al., 2024) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements. The security and privacy implications of such advanced sensing, from fingerprinting attacks (Huang et al., 2024b) to keystroke eavesdropping (Huang et al., 2024a), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data. Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (Liu et al., 2024b) and experience-driven model migration (Liu et al., 2024a), provide a roadmap for future privacy-preserving 3D capture ecosystems. This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects. We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (Wu et al., 2024).
@@ 2.1. Neural Scene Representations for Geometry
The foundation of modern neural rendering was laid by Neural Radiance Fields (NeRF) (Mildenhall et al., 2020), which represents a scene’s geometry and view-dependent appearance using a multilayer perceptron (MLP). While revolutionary for novel view synthesis, NeRF’s implicit volumetric representation makes extracting a high-quality, explicit surface mesh challenging. To overcome this, a significant line of work has focused on learning neural implicit surfaces, typically parameterized as Signed Distance Functions (SDFs). VolSDF (Yariv et al., 2021) and NeuS (Wang et al., 2021) were pioneering works in this direction, proposing novel formulations to convert SDFs into volume densities for rendering, thereby enabling the optimization of a coherent surface from 2D images. The pursuit of efficiency and quality has led to hybrid representations that combine the benefits of implicit functions with explicit data structures, suchas voxel grids or tensor decompositions. For instance, TensoIR (Jin et al., 2022) and K-Planes (Fridovich-Keil et al., 2023) utilize explicit low-rank tensor representations to achieve significantly faster training and rendering speeds. Other methods focus on ”baking” a trained neural field into an explicit structure like a mesh with features, enabling real-time rendering post-optimization (Chen et al., 2023a). The recent introduction of 3D Gaussian Splatting (Chen et al., 2022) has offered a powerful alternative, representing scenes as a collection of anisotropic Gaussians that can be rendered in real-time with high fidelity. The challenge of efficiently representing complex data is not unique to 3D vision; similar concerns motivate research in model compression (Fan et al., 2024) and the design of efficient architectures for tasks ranging from visual dialog (Guo et al., 2020) to crowd counting (Guo et al., 2019a).
@@ 2.2. Neural Inverse Rendering
Inverse rendering, the process of decomposing a scene into its fundamental components of geometry, material, and lighting, is a classic and notoriously ill-posed problem in computer vision. Neural approaches have made remarkable progress. Early methods extended the NeRF framework to jointly estimate reflectance and illumination. NeRFactor (Zhou et al., 2021) factorized the scene into a shape, a BRDF, and an illumination model represented by spherical harmonics. NeRD (Boss et al., 2021) further explored reflectance decomposition from challenging in-the-wild images. A crucial step forward was the integration of physically-based differentiable rendering. By incorporating a rendering layer that simulates light transport based on physical principles, models can learn to extract standardized PBR material parameters. PhySG (Zhang et al., 2021b) and IRON (Zhang et al., 2022b) were influential in demonstrating the joint optimization of neural SDFs with material properties like diffuse albedo and roughness.
The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (Munkberg et al., 2022), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images. Mat-NeuS (Zhang et al., 2022a) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties. Neural-PIL (Boss et al., 2022) introduced pre-integrated lighting to efficiently handle complex reflectance. These methods have been successfully applied to reconstruct general objects (Zhang et al., 2023) and even full indoor scenes (Yen-Chen et al., 2021; neil-net:2023). The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors. This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (Huang et al., 2020, 2023a) or building secure authentication systems resilient to fingerprinting attacks (Huang et al., 2024b). Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting. We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation (Guo et al., 2018, 2019b).
@@ 2.3. Generative Priors for 3D Content Creation
The rise of large-scale generative models, particularly 2D diffusion models, has unlocked new paradigms for 3D content creation. DreamFusion (Poole et al., 2023) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF. This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (Lin et al., 2023) improving resolution and efficiency. ProlificDreamer (Wang et al., 2023) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results. A key focus of recent work is the disentanglement of geometry and appearance within this generative framework. Fantasia3D (Chen et al., 2023c) and TexFusion (Seo et al., 2023) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts. This leverages the powerful multi-modal understanding of joint vision-language models like CLIP. Such models are also used for text-guided editing of existing 3D shapes (Richardson et al., 2023) or neural fields (Wang et al., 2022), enabling intuitive manipulation of scene properties (Dadashnialehi et al., 2023). The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog (Guo et al., 2021), video grounding (Li et al., 2021a), and even cross-modal emotion recognition (Gu et al., 2023b), all of which seek to bridge the gap between different data modalities.
@@ 2.4. Modeling of Advanced Materials and Dynamic Scenes
While significant progress has been made on opaque, Lambertian surfaces, modeling the full spectrum of real-world materials and dynamics remains an active research frontier. For complex materials, specialized representations are needed. NeRF-V (Zhang et al., 2021a) and NeRO (Srinivasan et al., 2023) explicitly model visibility and occlusion to handle complex shadowing and relighting. Polarized-NeRF (Mochida et al., 2022) leverages polarization cues to better separate diffuse and specular components. For translucent objects with subsurface scattering, methods like Neural Transmitted Radiance Fields (Yu et al., 2023) model the path of light through the object’s interior. Other works tackle anisotropic materials (Shen et al., 2023) or highly complex microstructures like hair (Chen et al., 2023b). The challenge of representing complex appearance is also addressed by works that learn decoupled neural textures (Baatz et al., 2022) or extract neural BRDFs (Wu et al., 2024).
Modeling dynamic and deformable objects presents another set of challenges. D-NeRF (Pumarola et al., 2021) pioneered the use of a deformation field to map a dynamic scene to a canonical space. This concept has been extended to create controllable and animatable avatars of human actors (Liu et al., 2021; Li et al., 2023), often incorporating relighting capabilities (Zhao et al., 2022). The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (Gu et al., 2023a). The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (Liu et al., 2022) or from distributed agents in a federated learning setting (Liu et al., 2024b, a). The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (Huang et al., 2024a). Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis (Zhou et al., 2022) or respiratory healthcare sensing (Zhao et al., 2025; Wang et al., 2024), is key.

### v2

@@ 1. Introduction
The demand for high-fidelity, interactive, and editable 3D content has become a cornerstone of modern digital experiences, fueling advancements in virtual and augmented reality, digital twins, and the entertainment industry . Traditionally, creating such photorealistic digital assets is a labor-intensive process, often requiring expert artists to manually model geometry and define complex material properties. The advent of neural scene representations has marked a paradigm shift, offering a new avenue for synthesizing and reconstructing the 3D world directly from images. Groundbreaking work on Neural Radiance Fields (NeRF) demonstrated an unprecedented ability to synthesize photorealistic novel views from a collection of images by representing a scene as a continuous volumetric function (mildenhall2020nerf). This seminal work spurred a wave of research exploring dynamic scenes (pumarola2021dnerf; li2021realtimedynamic) and even articulated subjects like humans (li2023tava; wu2023magicpony). However, the implicit volumetric nature of NeRF makes it challenging to extract explicit, high-quality surface geometry, which is crucial for most graphics pipelines.
@@
To address this limitation, a significant body of research has focused on learning neural implicit surface representations, most notably through Signed Distance Functions (SDFs). Frameworks like NeuS (wang2021neus) and VolSDF (yariv2021volsdf) ingeniously bridge the gap between volume rendering and surface rendering, enabling the extraction of detailed and coherent meshes from multi-view images. The pursuit of efficiency and detail in these reconstructions continues, with methods exploring explicit representations like tensor decompositions (jin2022tensoir; fridovich2023kplanes) and optimization ”baking” strategies to accelerate rendering (chen2023bakesdf). This hierarchical refinement of detail, from coarse volumetric fields to precise surfaces, mirrors the effectiveness of hierarchical models in other domains, such as understanding the nested structure of language in sign language translation or parsing complex scenes in visual dialog.
@@
Nevertheless, geometry alone is insufficient for true photorealism. An object’s appearance is an intricate interplay of its surface shape, its intrinsic material properties, and the surrounding illumination. The ultimate goal is thus inverse rendering: the challenging, ill-posed problem of disentangling these coupled factors from 2D observations. Early explorations in the neural domain sought to factorize NeRF into reflectance and illumination components (zhou2021nerfactor; boss2021nerd), often under simplified assumptions. Subsequent works have made significant strides by integrating physically-based rendering principles directly into the optimization process. These methods aim to decompose appearance into standardized Physically-Based Rendering (PBR) parameters—such as albedo, roughness, and metallic—by leveraging differentiable renderers (zhang2021physg; zhang2022iron). This disentanglement is critical, as failing to separate these components can lead to ambiguity, a challenge analogous to disambiguating events from correlated signals in audio-visual analysis or grounding language in proposal-free video understanding .
@@
State-of-the-art inverse rendering frameworks have achieved remarkable results, capable of extracting high-quality triangular meshes along with their PBR material maps and complex lighting environments (munkberg2022nvdiffrec; zhang2023invrender). These methods have been extended to handle complex indoor scenes (yen2021nerfilab; chang2023neilnet), unstructured ”in-the-wild” online image collections (liu2022neroic), and even animatable human avatars with relightable properties (zhao2022relightableavatar). The development of comprehensive frameworks has further unified these disparate efforts. Despite this progress, significant challenges remain. First, many high-fidelity methods are computationally expensive, hindering their practical application; recent advances in real-time relighting with 3D Gaussians (huang2024relightable3dgaussian) and efficient representations (chen2022splitgaussian) are beginning to address this. The need for efficient models is a universal concern, echoing the importance of model compression techniques like multi-objective convex quantization in deploying large networks . Second, most models are limited to simple, opaque materials, struggling with optically complex phenomena like translucency (yu2023neuraltransmitted), anisotropic reflectance (shen2023neraf), or the intricate structure of hair (chen2023neuralstrands). Third, robustly handling diverse real-world data requires sophisticated techniques, from modeling visibility and occlusion (zhang2021nerfv; srinivasan2023nero) to leveraging specialized information like polarization (mochida2022polarizednerf).
@@
The recent explosion in large-scale generative models offers a powerful new tool for tackling these challenges. Text-to-2D diffusion models have been ingeniously repurposed as powerful 3D priors through score distillation sampling (poole2022dreamfusion), enabling text-to-3D content creation (lin2023magic3d). These generative techniques are evolving to produce higher-fidelity and more diverse results (wang2023prolificdreamer) and are increasingly capable of disentangling geometry and appearance in a text-guided manner (chen2023fantasia3d). This allows for semantic texturing (richardson2023texture) and material assignment (seo2023texfusion), where textual concepts like ”metallic” or ”wooden” are translated into PBR values, a process informed by material-aware image generation (xu2023materialaware). Furthermore, multi-modal models like CLIP facilitate direct, text-driven editing of neural scenes (wang2022clipnerf), enabling intuitive control over object materials (dadashnialehi2023neuralscenedecoration) and textures (baatz2022nerftex). This convergence of vision, language, and graphics points towards a future of highly controllable and semantic 3D content creation, akin to how cross-modal understanding enhances tasks like sign language production and crowd counting .
@@
Beyond the visual spectrum, the broader field of sensing is also moving towards extracting nuanced information from ubiquitous signals. The ability to recognize human activities (huang2023phaseanti), gestures (gu2023attentionbased), emotions (gu2023wife), and even vital pulmonary functions (zhao2025wipulmo; wang2024targetoriented) from commodity WiFi signals showcases a parallel trend of disentangling complex information from noisy, indirect measurements. The security and privacy implications of such advanced sensing, from fingerprinting attacks (huang2024phyfinatt) to keystroke eavesdropping (huang2024keystrokesniffer), highlight a universal need for robust and secure systems, a consideration that will become paramount as 3D reconstruction moves towards distributed, user-sourced data. Architectures for such distributed systems, potentially leveraging federated learning with neural architecture search (liu2024finch) and experience-driven model migration (liu2024federated), provide a roadmap for future privacy-preserving 3D capture ecosystems. This thesis situates itself within this rich context, aiming to push the boundaries of high-fidelity, physically-based inverse rendering of general objects. We focus on robustly and efficiently disentangling geometry and complex material properties from multi-view images, drawing inspiration from both the explicit structure of physically-based rendering and the powerful priors of modern neural networks (wu2024extracting).
@@ 2.1. Neural Scene Representations for Geometry
The foundation of modern neural rendering was laid by Neural Radiance Fields (NeRF) (mildenhall2020nerf), which represents a scene’s geometry and view-dependent appearance using a multilayer perceptron (MLP). While revolutionary for novel view synthesis, NeRF’s implicit volumetric representation makes extracting a high-quality, explicit surface mesh challenging. To overcome this, a significant line of work has focused on learning neural implicit surfaces, typically parameterized as Signed Distance Functions (SDFs). VolSDF (yariv2021volsdf) and NeuS (wang2021neus) were pioneering works in this direction, proposing novel formulations to convert SDFs into volume densities for rendering, thereby enabling the optimization of a coherent surface from 2D images. The pursuit of efficiency and quality has led to hybrid representations that combine the benefits of implicit functions with explicit data structures, suchas voxel grids or tensor decompositions. For instance, TensoIR (jin2022tensoir) and K-Planes (fridovich2023kplanes) utilize explicit low-rank tensor representations to achieve significantly faster training and rendering speeds. Other methods focus on ”baking” a trained neural field into an explicit structure like a mesh with features, enabling real-time rendering post-optimization (chen2023bakesdf). The recent introduction of 3D Gaussian Splatting (chen2022splitgaussian) has offered a powerful alternative, representing scenes as a collection of anisotropic Gaussians that can be rendered in real-time with high fidelity. The challenge of efficiently representing complex data is not unique to 3D vision; similar concerns motivate research in model compression and the design of efficient architectures for tasks ranging from visual dialog to crowd counting .
@@ 2.2. Neural Inverse Rendering
Inverse rendering, the process of decomposing a scene into its fundamental components of geometry, material, and lighting, is a classic and notoriously ill-posed problem in computer vision. Neural approaches have made remarkable progress. Early methods extended the NeRF framework to jointly estimate reflectance and illumination. NeRFactor (zhou2021nerfactor) factorized the scene into a shape, a BRDF, and an illumination model represented by spherical harmonics. NeRD (boss2021nerd) further explored reflectance decomposition from challenging in-the-wild images. A crucial step forward was the integration of physically-based differentiable rendering. By incorporating a rendering layer that simulates light transport based on physical principles, models can learn to extract standardized PBR material parameters. PhySG (zhang2021physg) and IRON (zhang2022iron) were influential in demonstrating the joint optimization of neural SDFs with material properties like diffuse albedo and roughness.
The current state-of-the-art is represented by powerful frameworks like NVDIFFREC (munkberg2022nvdiffrec), which directly optimizes a triangle mesh, its PBR materials, and an HDR light map to match input images. Mat-NeuS (zhang2022matneus) extends the high-fidelity surface reconstruction of NeuS to simultaneously acquire material properties. Neural-PIL (boss2022neuralpil) introduced pre-integrated lighting to efficiently handle complex reflectance. These methods have been successfully applied to reconstruct general objects (zhang2023invrender) and even full indoor scenes (yen2021nerfilab). The core challenge in all these methods is robust disentanglement, which requires careful regularization and priors. This is conceptually similar to the need for robust feature extraction in other domains, such as developing anti-interference techniques for WiFi-based activity recognition (huang2023phaseanti) or building secure authentication systems resilient to fingerprinting attacks (huang2024phyfinatt). Our work contributes to this line of research by proposing new techniques to improve the robustness of material decomposition under unknown lighting. We also draw inspiration from how hierarchical models can parse complex structures, a principle proven effective in tasks like sign language translation.
@@ 2.3. Generative Priors for 3D Content Creation
The rise of large-scale generative models, particularly 2D diffusion models, has unlocked new paradigms for 3D content creation. DreamFusion (poole2022dreamfusion) introduced Score Distillation Sampling (SDS), a method to ”distill” the knowledge of a pre-trained 2D text-to-image model to guide the optimization of a 3D representation like NeRF. This spawned a flurry of research in text-to-3D generation, with follow-up works like Magic3D (lin2023magic3d) improving resolution and efficiency. ProlificDreamer (wang2023prolificdreamer) proposed Variational Score Distillation (VSD) to enhance the diversity and fidelity of generated results. A key focus of recent work is the disentanglement of geometry and appearance within this generative framework. Fantasia3D (chen2023fantasia3d) and TexFusion (seo2023texfusion) explicitly aim to generate not just a shape, but also assign meaningful textures and PBR materials based on textual prompts. This leverages the powerful multi-modal understanding of joint vision-language models like CLIP. Such models are also used for text-guided editing of existing 3D shapes (richardson2023texture) or neural fields (wang2022clipnerf), enabling intuitive manipulation of scene properties (dadashnialehi2023neuralscenedecoration). The ability to generate content based on abstract, multi-modal inputs resonates with research in visual dialog , video grounding , and even cross-modal emotion recognition (gu2023wife), all of which seek to bridge the gap between different data modalities.
@@ 2.4. Modeling of Advanced Materials and Dynamic Scenes
While significant progress has been made on opaque, Lambertian surfaces, modeling the full spectrum of real-world materials and dynamics remains an active research frontier. For complex materials, specialized representations are needed. NeRF-V (zhang2021nerfv) and NeRO (srinivasan2023nero) explicitly model visibility and occlusion to handle complex shadowing and relighting. Polarized-NeRF (mochida2022polarizednerf) leverages polarization cues to better separate diffuse and specular components. For translucent objects with subsurface scattering, methods like Neural Transmitted Radiance Fields (yu2023neuraltransmitted) model the path of light through the object’s interior. Other works tackle anisotropic materials (shen2023neraf) or highly complex microstructures like hair (chen2023neuralstrands). The challenge of representing complex appearance is also addressed by works that learn decoupled neural textures (baatz2022nerftex) or extract neural BRDFs (wu2024extracting).
Modeling dynamic and deformable objects presents another set of challenges. D-NeRF (pumarola2021dnerf) pioneered the use of a deformation field to map a dynamic scene to a canonical space. This concept has been extended to create controllable and animatable avatars of human actors , often incorporating relighting capabilities (zhao2022relightableavatar). The ability to track and model human motion finds parallels in other sensing modalities, such as gesture recognition using WiFi (gu2023attentionbased). The overarching goal is to build robust models from diverse, often unstructured data, whether it’s from online image collections (liu2022neroic) or from distributed agents in a federated learning setting (liu2024finch; liu2024federated). The security and privacy of such data are paramount, a concern highlighted by research into the vulnerabilities of sensing systems (huang2024keystrokesniffer). Our work focuses on static objects but aims for a material representation robust enough to serve as a foundation for future extensions into these more complex dynamic and material domains, where understanding subtle signals, much like in audio-visual event analysis or respiratory healthcare sensing (zhao2025wipulmo; wang2024targetoriented), is key.